OpenStudy (anonymous):

find arc of a circle when r=13.9 and theta: 150 degrees

OpenStudy (anonymous):

[150 /180 ]* 3.14 = 2.62 radian ==> arc length = 2.62 * 13.9 =36.42

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

5*pi/6 *13.9=139/12*pi 36.39